OpenStudy (anonymous):

f(-2)=|-6x^(2)-x^(4)+4x|

OpenStudy (bahrom7893):

okay what are we looking for?

myininaya (myininaya):

are you suppose to plu in the -2 where the x's are?

myininaya (myininaya):

f(x)=|-6x^2-x^4+4x| and you want to find f(-2) right?

OpenStudy (saifoo.khan):

plug*

OpenStudy (saifoo.khan):

lol.

OpenStudy (bahrom7893):

Ohh lol I see

OpenStudy (anonymous):

it says find f(2)

myininaya (myininaya):

lol shhh khan

OpenStudy (saifoo.khan):

hehe! ;)

OpenStudy (bahrom7893):

f(2) or f(-2)?

OpenStudy (anonymous):

f(-2)

myininaya (myininaya):

f(-2)=|-6(-2)^2-(-2)^4+4(-2)|

OpenStudy (saifoo.khan):

She got it.

OpenStudy (saifoo.khan):

i mean, Myinin got it right.

OpenStudy (anonymous):

f(x)=|-6x^(2)-x^(4)+4x| find f(-2) That was the directions

OpenStudy (saifoo.khan):

then u have to do this: f(-2)=|-6(-2)^2-(-2)^4+4(-2)|

myininaya (myininaya):

use your order of operations

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

would there be two answers then because of the absolute value signs

myininaya (myininaya):

no

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

hello?

myininaya (myininaya):

|-3| means we want the distance from -3 to 0 which is 3 units right? |-3|=3 it only gives one distance

myininaya (myininaya):

i was typing lol you are impatient

OpenStudy (saifoo.khan):

but if u hav: |x| = 3 then the values be 3 or -3

OpenStudy (anonymous):

oh sorry lol

myininaya (myininaya):

yes saifoo is exacly right

OpenStudy (saifoo.khan):

:D

OpenStudy (saifoo.khan):

thnks!

OpenStudy (anonymous):

yes i that is what i thought

OpenStudy (anonymous):

so would the answer be -64 and 64?

myininaya (myininaya):

no you only get one answer

OpenStudy (saifoo.khan):

there is a difference: |x| = 3 then the values be 3 or -3 But, if u hav x= |-3| then x = 3

OpenStudy (anonymous):

ooohhh ok i get it

OpenStudy (anonymous):

so the answer is 64 correct?

myininaya (myininaya):

i'm getting 48 not 64

OpenStudy (anonymous):

f(-2) = |-6(-2)^2 - (-2)^4 + 4(-2)| f(-2) = |-6(4) - (32) - 8| f(-2) = |-24 - 32 - 8| f(-2) = |-64| f(-2) = 64

myininaya (myininaya):

(-2)^4=?

OpenStudy (saifoo.khan):

16

myininaya (myininaya):

(-2)^4=(-2)(-2)(-2)(-2)=4(4)=16

OpenStudy (anonymous):

oh

myininaya (myininaya):

so f(-2)=|-6(-2)^2-(-2)^4+4(-2)| =|-6(4)-16-8| =|-24-16-8|

OpenStudy (anonymous):

it's -48

myininaya (myininaya):

=|-48|=48

OpenStudy (saifoo.khan):

lol.

OpenStudy (anonymous):

oh great!

OpenStudy (anonymous):

Thanks !

myininaya (myininaya):

|-48| is telling us to find the distance from -48 and 0 and distance is zero or positive so the distance from -48 to 0 is 48 and there not two answers because you can only have one number to represent the distance

OpenStudy (anonymous):

oh ya that make sense thank you so much!

myininaya (myininaya):

np

OpenStudy (saifoo.khan):

myinin, i just saw.. v both are going with same Pace! ;)

OpenStudy (saifoo.khan):

Same Level, same # of Medals!

OpenStudy (anonymous):

you both are awesome lol!

OpenStudy (saifoo.khan):

lol.